If the moon has a core with a radius of 400km and the uniform density of the overlying mantle is 3300kgm-3, what is the density of the core, hence calculate the mass of the core.


1
Expert's answer
2021-03-18T15:23:48-0400

The radius of the Moon is "1738\ km" and its mean density is "3347 \ kg/m^3" . So,


mass of core "m_c= \frac{4}{3}\pi r_c^3ρ_c"


volume of mantle "V=\frac{4}{3}\pi R^3−\frac{4}{3}\pi r_c^3"


mass of mantle "m_m= \frac{4}{3}\pi (R^3-r_c^3)\rho_m"


mass of planet "m=m_m+m_c"


"\frac{4}{3}\pi R^3\rho= \frac{4}{3}\pi (R^3-r_c^3)\rho_m+ \frac{4}{3}\pi r_c^3ρ_c\to"


"\frac{\rho-\rho_m}{\rho_c-\rho_m}=(\frac{r_c}{R})^3\to\frac{3347-3300}{\rho_c-3300}=(\frac{400}{1738})^3\to"


"\rho_c\approx7155\ kg/m^3" . Answer